4. What is the probability of having the 64th head on the 99th toss?
1
Expert's answer
2012-01-31T11:27:05-0500
We can arrange 63 heads on 98 tosses in n = (98,63) = 98!/(63!(98-63)!) = 460149430026457132088731968 ways.
The total number of possible outcomes is N = 2^98 = 316912650057057350374175801344.
The probability of having 63 heads in 98 tossings is p = n/N = 460149430026457132088731968/316912650057057350374175801344 ≈ 0.0014519755836304144138101871749498.
So, the probability the 64th head on the 99th toss is P = p*(1/2) ≈ 0.0007259877918152072069050935874749.